TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Create new blog entry for Point 4


Created by Main.BorisBellesia on 2007-11-20 - 19:26

Shift HC teams and Piquet numbers - 20th Nov. 2007

PDF shifts.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x20x19x25/shifts.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-19 - 20:48

RAT CCC: Powering tests of Sector45 - 19 Nov. 2007

PDF RAT_20071119.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x19x20x47/RAT_20071119.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-14 - 18:30

RAT CCC: preparation for Powering test: Sector45 - 14th Nov.

PDF RAT_20071114.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x14x18x29/RAT_20071114.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-11-13 - 19:09

RAT Point4 - Preparation for Powering Sector45 - 9th Nov. 2007

PDF RAT_20071112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x11x13x19x07/RAT_20071112.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-31 - 10:40

From Pressure Test to Cooldown - Sectors 45 and 34 - 30th October

Pptx Sector45_34_status_20071030.pptx:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x31x10x10/Sector45_34_status_20071030.pptx

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-23 - 12:24

From pressure test to Cooldown - Sector45 and 34 - 23rd October

Pptx Sector45_34_status_20071023.pptx:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x23x12x22/Sector45_34_status_20071023.pptx

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-23 - 12:21

From pressure test to cooldown - 23rd October 2007

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-23 - 12:15

Preparation for cooldown - RF-Powering tests-Installation activities

PDF RAT_20071022.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x23x12x14/RAT_20071022.pdf

PowerPoint Sector45_34_RF-PT_planning.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x23x12x14/Sector45_34_RF-PT_planning.ppt

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-16 - 16:06

From Pressure test to Cooldown - 16th October - Point4

Pptx Sector45_34_status_20071015.pptx:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x16x15x57/Sector45_34_status_20071015.pptx

Sectors 45 and 34

  • Minutes: see attached presentation

  • Sector34 Open Issue: Needed strict coordination among AT-MCS, AT-VAC and AT-ACR to define priorities of DSL interconnection and 7L3 leaks localization/repair. A. Perin is the expert to who address technical questions.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-16 - 14:24

From Pressure test to CooLdown - 16th Oct.

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-10-10 - 20:21

Warm Magnets Commissioning Point 4. Minutes 09-10-2007

PDF Warm_Magnets_Point4_Minutes.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x10x10x20x20/Warm_Magnets_Point4_Minutes.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-10-09 - 20:16

From Pressure Test to Cool-Down - Tuesday 09 - Point 4

SECTOR45

General planning:

  • W41-W42: L5 to 80K
  • W42-W43: Cool-down of the whole sector from 80K up to 4.5K
  • W44: Cool-down of the arc45 from 4.5K up to 1.9K / ElQA of the Circuits of the matching sections / Powering tests of the matching sections
  • W45: ElQA of the circuits of the arc and Powering tests of the matching sections
  • W46: Powering tests of the circuits of the arc

Activitites - w41:

  • Today -09/10 - cooldown L5 to 80K
    • Sector closed since 08/10
    • Access allowed with key and tokens except RE52-UL55
    • See document LHC-S-ES-0017 for particular access conditions
  • ElQA break-down monitoring all along cooldown; AT-MEL
  • QPS activities; AT-MEL
  • CMS vacuum installation/tests; AT-VAC
  • Magnet alignment; TS-SU
  • Cable campaign point5; TS-EL
  • Inner triplet interconnection; AT-MCS

Open Issues + Long term activitites:

  • Swapped cables of RB circuit on DFBA RR53
  • Cleaning – W43/44
  • Collimator installation W44
  • CMS-ZDC optical fiber installation; W44

SECTOR34

General planning+Activities:

  • W40: Pressure test of the arc / DSL still disconnected
  • W41-W42:
    • Consolidation
      • All lines under nominal pressure up to Wed. 10/10 - AT-VAC/ACR
      • Cryo-valves reconfiguration Thu.-Fri. 11-12/10 - AT-ACR
      • Opening and repairs of leaks (two leaks found: 7R4 ~2*10^-6 [mbar*l/s] and 7L4 ~10^-5 [mbar*l/s]) - AT-MCS/VAC
    • Checks the jacks - stable pressure conditions needed - 2 days needed - date to be decided, possibly on Friday 12/10 - TS-IC
    • ElQA.DOC from Wed. 10/10 - AT-MEL
    • Manometer installation on the ceiling at the level of RE38 – trolley passage difficult – TS-IC
  • W43: DFBLC-DSLC-DFBAF interconnections start (possibility to start during W42 if compatibility with consolidation)
  • W50: Flushing

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-09-25 - 16:44

Sector45 - Access conditions and cool-down co-activities

PDF Access_restart_cd_20070925.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x09x25x16x43/Access_restart_cd_20070925.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-29 - 16:33

Activities foreseen in sector 45 from Monday 2nd July up to the cooldown

Activities in sector 45 from Monday 2nd July

  • CRYO
    • Purging on Friday 29th evening and all the week end long;
    • Refilling on Monday 2nd July;
  • ElQA-TP4B and Monitoring installation
    • On Tuesday 3th July, once 6Bar conditions are reached, ElQA will be repeated (only on RB) maybe from 8am to 1pm. RESTRICTED ACCESS.
  • CRYO-instrumentation
    • Coherence test of all the valves in ARC and LSS (Thursday 28th and Friday 29th);
    • Commissioning;
    • Initialization of the CV910 valves on Tuesday 26th and Wednesday 27th;
  • TS/EL intervention
    • Changing of the connectors of a BPM (LSS4R);
    • Connection check of the instrumentation cables of BCT;
  • TIM camera
    • General check on Monday 2nd July;
    • Run from P4 to P5, during the Cryo Patrouille the morning of the start of the cool-down;
  • Safety Inspection
    • Monday 2nd July at 2pm;
  • AT/ACR Patrouille
    • A Patrouille will be carried out on Wednesday 4th July just before the cooldown.
  • AT/VAC
    • Short access to LSS4R and LSS5L on Monday 2nd July.
  • AB/BI
    • LSS4R short intervention on the BI (capeau).

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-27 - 18:52

going on activities in sector 45

Going on activities in sector 45

  • DFBAH and MB (diode)
    • checked and closed by Wednesday 27th;
    • Leak test on Thursday 28th;
    • Purging on Friday 29th evening and all the week end long;
    • Refilling on Monday 2nd July;
    • On Tuesday 3th July, once 6Bar conditions are reached, ElQA will be repeated (only on RB);
  • CRYO-instrumentation
    • Initialization of RF valves on Wednesday 27th or Thursday 28th;
    • Coherence test of all the valves in ARC and LSS (Thursday 28th and Friday 29th);
    • Commissioning;
    • Initialization of the CV910 valves on Tuesday 26th and Wednesday 27th;
  • TS/EL intervention
    • Changing of the connectors of a BPM (LSS4R);
    • Placing of some Radmon cables on Wednesday 27th, Thursday 28th and Friday 29th (LSS5L);
    • Connection check of the instrumentation cables of BCT;
  • TIM camera
    • General check on Monday 2nd July;
    • Run from P4 to P5, during the Cryo Patrouille the morning of the start of the cool-down;
  • Safety Inspection
    • Monday 2nd July at 2pm;
  • RAMSES
    • Check of the equipment on Tuesday 26th and Wednesday 27th.
  • CV Intervention
    • Thursday 28th June from 2pm to 4pm
  • QUI
    • Caps removal, sleeve installation and then closure of the QUI performed on Thursday 28th.
    • The new mechanism for the header B valve is being prepared and should be ready for Friday 29th or early next week.

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2007-06-26 - 18:47

Diagnosis of MB.A45 circuit shorted to ground - Courtesy of D.Bozzini & J.P.Tock

This morning MCS has open the diode box of the dipole 3186. AT/MEL has connected an ohmmeter during the operation to follow the status of the short. At the time of opening the big flange of the diode box the resistance to ground started to change (increase). After the opening it could be possible to first see a quite big quantity of dust on the diode metallic box and second some big pieces of welding residuals.

After having disconnected the diode at the level of the half moons the short disappeared (infinite resistance) and in at the same time some debris felled down from the top part of the half moons. As you can see in the attached pictures the size of one of this residual fits perfectly between the live part and the metallic body of the bus bar duct. In addition in one picture you will see a carbon trace on the G10 insulating piece. This could be provoked by the current of 1A we injected to better localize the fault position flowing through the debris supported onto the G10 piece.

The system will be set back to the requested pressure condition (6 bars in the magnets, 2.2 bars in the DFBA's - which should be achieved by Tuesday 3rd July), when the ELQA-TP4B qualification of this faulty circuit will be repeated.

JPEG IMG_0871.jpg:
IMG_0871.jpg

JPEG IMG_0933.jpg:
IMG_0933.jpg

JPEG IMG_0909.jpg:
IMG_0909.jpg

JPEG IMG_0882.jpg:
IMG_0882.jpg

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2007-06-13 - 18:14

Beam Instrumentation - BSRT Installed and Aligned

The two BSRT telescopes were Installed last Monday in LSS4.

LSS4L: The BSRTD has been connected to the BSRTM and the system successfully aligned with the laser.

LSS4R: The vacuum chamber in IP4 right side is still intact after the BGI bake-out incident and therefore connection and alignment of the BSRT right side will be done tomorrow !

J.J. Gras for the BI Team

JPEG BSRTD_C5L4_Front1.jpg:
BSRTD_C5L4_Front1.jpg

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-08 - 18:42

RAT in Point 4 - Minutes of 8th June 2007

PDF RAT_20070608.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x06x08x18x41/RAT_20070608.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2007-06-04 - 00:37

RAT in P4 - Minutes 1st June 2007 (3 weeks for Cool-Down !)

PDF RAT_20070601.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x06x04x00x35/RAT_20070601.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-22 - 18:34

RAT in Point 4 - Minutes of 22nd May 2007

PDF RAT_20070522.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x05x22x18x33/RAT_20070522.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-22 - 18:33

RAT in Point 4 - Minutes of 11th May 2007

PDF RAT_20070511.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x05x22x18x32/RAT_20070511.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-07 - 18:38

RAT in Point 4 - Minutes of 4th May 2007

PDF RAT_20070504.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x05x07x18x37/RAT_20070504.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-04-27 - 17:47

RAT in Point 4 - Minutes of 27th April 2007

PDF RAT_20070427.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x04x27x17x39/RAT_20070427.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-04-20 - 18:36

RAT in Point 4 - Minutes of 20th April 2007

PDF RAT_20070420.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x04x20x18x26/RAT_20070420.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-04-17 - 18:52

RAT in Point 4 - Minutes of 13th April 2007

PDF RAT_20070413.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x04x17x18x51/RAT_20070413.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-03-30 - 18:47

RAT in Point 4 - Minutes of 30th March 2007

PDF RAT_20070330.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x03x30x18x45/RAT_20070330.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-03-28 - 09:50

RAT in Point 4 - Minutes of 23th March 2007

PDF RAT_20070323.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x03x28x09x49/RAT_20070323.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-03-21 - 10:56

RAT in point 4 - Minutes of 16th March

PDF RAT_20070316.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x03x21x10x54/RAT_20070316.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-03-12 - 18:57

RAT in Point 4 - Minutes of 9th March 2007

PDF RAT_20070309.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x03x12x18x56/RAT_20070309.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-02-23 - 17:03

RAT in point 4 - minutes of the 23th February 2007

PDF RAT_20070223.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x02x23x17x01/RAT_20070223.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-02-16 - 18:46

RAT in point 4 - minutes of 16th February 2007

PDF RAT_20070216.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x02x16x18x45/RAT_20070216.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-09-25 - 17:30

Dry-Run RAT 06-09-21

PDF Z:\Antonio\Dry-Run\RAT\DR_RAT_060921.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x09x25x17x28/Z:\Antonio\Dry-Run\RAT\DR_RAT_060921.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-30 - 10:33

24-h run UA43: RATs first round table

Minutes of the meeting held this morning as usual 8h30 in CCC.

First round table related to the 24 hours heat run in UA43 and UJ43.

Have a nice day

PDF RAT_20060830.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x30x10x31/RAT_20060830.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-29 - 16:04

RATs: 24-h heat run completed today 29.08 at 15h20

Last news from the field (UA43 and UJ43)

24-h run completed today at 15h20

Please find enclose few more details concerning these last two days of HC activity in point 4.

Regards

PDF RAT_20060829.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x29x16x01/RAT_20060829.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-28 - 11:17

RAT 24-hours test today

Minutes of the last RAT meeting

Today 24-h heat run in UA43 from 14h00

PDF RAT_20060828.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x28x11x14/RAT_20060828.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-25 - 11:07

RAT: first analysis of the 8h-run and RR53

Minutes of the meeting, including:

1) First analysis related to the 8h-run (24.08 from 9h27 to 17h05)

2) Round table concerning the work advancement in RR53

PDF RAT_20060825.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x25x11x04/RAT_20060825.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-24 - 11:03

RAT: 8 hours run today..

Minutes from the last meeting in CCC, mainly related to the 8 hours run (in progress)..

PDF RAT_20060824.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x24x11x01/RAT_20060824.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-23 - 11:29

Last from RATs...

Please find enclose the minutes of the last RATs meeting, held in point 4 this morning 8h30, for the preparation of the LHC short-circuit test in UA43.

PDF RAT_20060823.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x23x11x27/RAT_20060823.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-21 - 10:34

News from tunnel RATs

Here you are fresh news from tunnel RATs.

Enclose the minutes of last RAT meeting, held today in SX4.

Infos related to the interlock tests and the next 8h-run.

Regards

PDF RAT_20060821.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x21x10x28/RAT_20060821.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-17 - 11:18

RAT 4you

Minutes of last RAT meeting this morning in point 4. Main issues: next 8h run and 24h run end of month.

PDF RAT_20060817.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x17x11x16/RAT_20060817.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-15 - 10:59

Last RAT news: interlock test RQD and RQF

Please find enclose the minutes of the last RAT meeting, held in point 4 this morning. The doc includes infos related to the last interlock test and the future 8hrs-run.

PDF RAT_20060815.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x15x10x57/RAT_20060815.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-08-14 - 11:00

RAT meeting today..

Please find enclose the last RAT meeting minutes.. 1st 8h run in UA43, interlock test for RQD and RQF, and next steps.

PDF RAT_20060814.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2006x08x14x10x57/RAT_20060814.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2006-07-17 - 15:25

des tests HTdes cables refroidis de UA43

Les tests d isolement des cables refroidis de l UA43 ne s averent pas tous positifs... Apres bien des interrogations , il semble que l eau contenue dans les cables soit de tres mauvaise qualite et perturbe considerablement les mesures ... Apres discutions avec Loic Bougeois, il ne semble pas d actualite de vidanger ces cables refoidis mais plutot de faire circuler l eau demineralisee en vue de la debarrasser des impuretes qu elle contient... Travail effectue apres la pose des eletas...et le raccordement aux switchs.

JPEG sertissage__trop__pres___RA43.jpg:
sertissage__trop__pres___RA43.jpg

JPEG sertissage_trop_pres__RA43.jpg:
sertissage_trop_pres__RA43.jpg

Comments:

  • et puis nous avons finalement vidange les cables et tout est rentre dans l ordre le jour suivant....les resistances d isolement sont conformes a ce qu on attendait... -- Main.JeanPierreMalodDognin - 21 Jul 2006 - 16:43

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-04-06 - 15:17

proche des masques Africains?

un detail de DQS , pour rever un peu...

Comments:

JPEG DQSdetail.JPG:
DQSdetail.JPG

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:53

RAT meeting minutes 3rd of March 2006

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:50

RAT meeting minutes 2nd of March 2006

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:48

RAT meeting minutes 1st of March 2006

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:49

RAT meeting minutes 27th of February 2006

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-03-30 - 08:46

RAT meeting minutes 24th of February 2006

Comments:

Edit, attach file or add comment


Created by Main.RobertoSaban on 2006-03-30 - 08:36

Field Control Room in UA43

At the request of Sebastien Pelletier from TS-MEL/Cabling Section, I asked Pierre Charrue and his team to remove all the computing equipment in UA43. The PCs, monitors and the big screen are now stored by AB-CO. A request to re-install this must be made to them a month in advance of the start of the Short Circuit Tests in UA43 en of July 2006.

Comments:

-- Main.JeanPierreMalodDognin - 05 Apr 2006

Et, hier 4 avril...pour en rajouter une couche(!) dans le demontage(provisoire?) de le SdeC,j ai -apres demande de Rosario- demonte les baches qui en cachait l interieur ...la video surveillance est toujours operationnelle et l acces par code conserve.

-- Main.JeanPierreMalodDognin - 05 Apr 2006

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-01-09 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login